Вы находитесь на странице: 1из 102

Indian Oiltanking Limited

(BOOT-3, South Oil Jetty)


Paradip Refinery, Paradip

Date:01-10-2019
Limited
l Jetty)
Paradip

2019
Indian Oiltanking Limited
(BOOT-3, South Oil Jetty)
Paradip Refinery, Paradip

Date:01-10-2019
TOWER MONITOR (At Jetty Top)

MOV STATUS
HORIZONTAL VERTICAL JET/FOG FOAM LEVEL IN
Tower Monitor No
MOVEMENT STATUS MOVEMENT STATUS MOVEMENT STATUS FOAM TANK - 1
WATER FOAM

TM - 1 OK OK OK OK Not OK

TM - 2 OK OK OK OK OK
126cm out of 190
cm
TM - 3 OK OK OK OK Not OK

TM - 4 OK OK OK OK Not OK

SELF OSCILLATING / BASE MONITOR (At Jetty Top)

FOAM LEVEL
HORIZONTAL VERTICAL JET/FOG OPERATING FOAM LEVEL IN
Base Monitor No IN FOAM
MOVEMENT STATUS MOVEMENT STATUS MOVEMENT STATUS HOV STATUS FOAM TANK - 2
TANK - 1

BM - 1 OK OK OK OK

126 cm out of 118cm out of 190


190 cm cm
BM - 2 OK OK OK OK
126 cm out of 118cm out of 190
190 cm cm
BM - 3 OK OK OK OK

BM - 4 OK OK OK OK

QUICK RELEASE MOORING HOOK (At Jetty Top)


Hook-1 Release Operation Hook-1 Lock Hook-2 Release Operation Hook-2 Lock
QRMH No
Electrically Mechanically Operation Electrically Mechanically Operation

Mooring Dolphin - 1 OK OK OK OK OK OK

Mooring Dolphin - 2 OK OK OK OK OK OK

Berthing Dolphin - 1 OK OK OK OK OK OK

Berthing Dolphin - 2 OK OK OK OK OK OK

Mooring Dolphin - 3 OK OK OK OK OK OK

Mooring Dolphin - 4 OK OK OK OK OK OK

Diesel Engine Driven Fire Water Pump (At FWPH)


Fire Water Engine No Running Condition Last Diesel Level Last Run on Any other Observation

90-PA-103 OK 72.34% 30.09.2019

90-PA-104 OK 74.67% 22.09.19


90-PA-105 OK 34.99% 30.09.2019

90-PA-106 OK 58.76% 30.09.2019

90-PA-107 OK 63.98% 30.09.2019

90-PA-108 OK 98.66% 30.09.2019 LT breakdown

Electric Driven Fire Water Jockey Pump (At FWPH)


Fire Water Engine No Running Condition Last Run On Any other Observation Remarks
(Auto/Manual)
90-PA-101A Manual 7/28/2019 Feeder damage

90-PA-101B Manual 27.09.2019

90-PA-102A Manual 8/12/2019 Abnormal sound observed during running

90-PA-102B Manual 27.09.2019


mited
y)
dip

FUNCTIONALITY OF
LAST
FOAM LEVEL IN INLINE BALANCE
CHECKED REMARKS
FOAM TANK - 2 PRESSURE
DATE
PROPORTIONER

30.09.2019 Foam Line MOV not


serviceable.

30.09.2019
118cm out of 190cm Foam line filled
30.09.2019 Foam Line MOV not
serviceable.

Foam Line MOV not


30.09.2019
serviceable.

Top)
FUNCTIONALITY
OF INLINE FUNCTIONALITY OF LAST
BALANCE SELF OSCILLATING CHECKED REMARKS
PRESSURE GEAR SYSTEM DATE
PROPORTIONER

Not operating since 29.09.2019


beginning

Foam line filled


Not operating since
beginning 29.09.2019

Foam line filled


Not operating since
beginning 29.09.2019

Not operating since


29.09.2019
beginning

Top)
Capstan Operation Last Checked
Remarks
By Pedestal By Switch Date

OK OK 30.09.2019

OK OK 30.09.2019

OK OK 30.09.2019

OK OK 30.09.2019

OK OK 30.09.2019

OK OK 30.09.2019

H)
Remarks
`
Indian Oiltanking Limited
(BOOT-3, South Oil Jetty)
Paradip Refinery, Paradip
Date:01-10-2019
OPERATION
DESCRIPTION OF THE ACTIVITY JOB STATUS DONE BY REMARKS

Fire engines daily inspection done. Completed Darshan P Checklist updated


Fire Engine-103 & 105 trial runs taken in run mode for 30 min each. Completed Darshan P Checklist updated
Fire Engine-106, 107 & 108 trial runs taken in idle mode for 05 min each. Completed Darshan P Checklist updated
MECHANICAL
DESCRIPTION OF THE ACTIVITY JOB STATUS DONE BY REMARKS

Maintenance activity of XZV-0026 Completed Biplab M


Repairing /Chipping/Bio-guard application on process pipeline Under progress Biplab M
ELECTRICAL ACTIVITIES
DESCRIPTION OF THE ACTIVITY JOB STATUS DONE BY REMARKS

PM of ESD battery bank Under progress Pratap P


Healthiness checking of DG battery charger Under progress Pratap P
INSTRUMENTATION
DESCRIPTION OF THE ACTIVITY JOB STATUS DONE BY REMARKS
----------- --------- ----------- ------------
JOB BY OTHER AGENCY
DESCRIPTION OF THE ACTIVITY JOB STATUS DONE BY REMARKS

Cable lying and panel board installation at south side of control room Under progress B/R Co.

PERMIT DETAILS SOJ AREA Date:01-10-2019


PERMIT No
Cold 2
Hot 0
Vehicle 0
Confined space 0
Work at height 0
Excavation work 0
Electrical / Energisation-De-energisation 2
Radiation work 0
TOTAL 4

S
SL REQUES
DATE
No. T No.

1 01.04.2018 1

2 02.04.2018 2

3 03.04.2018 3

4 04.04.2018 4

5 04.04.2018 5

6 05.4.2018 6

7 08.04.2018 7

8 09.04.2018 8

9 10.04.2018 9

10 10.04.2018 10

11 10.04.2018 11

12 11.04.2018 12

13 11.04.2018 13

14 13.04.2018 14
15 13.04.2018 15

16 14.04.2018 16

17 16.04.2018 17

18 17.04.2018 18

19 20.04.208 19

20 22.04.2018 20

21 25.04.2018 21

22 03.05.2018 22

23 05.05.2018 23

24 07.05.2018 24

25 07.05.2018 25

26 10.05.18 26

27 11.05.18 27

28 12.05.18 28

29 12.05.18 29

30 12.05.18 30

31 15.05.18 31

32 17.05.18 32

33 18.05.18 33

34 18.05.18 34
35 19.05.18 35

36 21.05.18 36

37 21.05.18 37

38 22.05.18 38

39 23.05.18 39

40 24.05.18 40

41 25.05.18 41

42 25.05.18 42

43 29.05.18 43

44 29.05.18 44

45 29.05.18 45

46 30.05.18 46

47 31.05.18 47

48 31.05.18 48

49 31.05.18 49

50 31.05.18 50

51 31.05.18 51

52 31.05.18 52

53 03.05.18 53

54 05.06.18 54

55 05.06.18 55

56 05.06.18 56
57 09.06.2018 57

58 10.06.18 58

59 10.06.18 59

60 10.06.18 60

61 10.06.18 61

62 11.06.18 62

63 11.06.18 63

64 13.06.18 64

65 13.06.18 65

66 14.06.18 66

67 14.06.18 67

68 15.06.18 68

69 16.06.18 69

70 17.06.18 70

71 17.06.18 71

72 17.06.18 72

73 20.06.18 73

74 22.06.18 74

75 22.06.18 75

76 22.06.19 76

77 22.06.20 77

78 23.06.2018 78
79 23.06.2018 79

80 23.06.2018 80

81 24.06.18 81

82 24.06.18 82

83 25.06.18 83

84 26.06.18 84

85 27.06.2018 85

86 28.06.2018 86

87 28.06.2018 87

88 01.07.2018 88

89 01.07.2018 89

90 01.07.2018 90

91 01.07.2018 91

92 03.07.2019 92

93 04.07.2018 93

94 05.07.2018 94

95 05.07.2018 95

96 08.07.18 96

97 08.07.18 97

98 09.07.18 98

99 11.07.18 99

100 11.07.18 100

101 11.07.18 101

102 11.07.18 102

103 11.07.18 103


104 06.07.18 104

105 06.07.18 105

106 06.07.18 106

107 17.07.18 107

108 17.07.18 108

109 21.07.18 109

110 24.07.18 110

111 25.07.18 111

112 25.07.18 112

113 25.07.18 113

114 26.07.18 114

115 27.07.18 115

116 30.07.18 116

117 06.08.18 117

118 11.08.18 118

119 13.08.187 119

120 16.08.2018 120

121 20.08.2018 121

122 21.08.2018 122

123 21.08.2018 123

124 21.08.2018 124

125 23.08.2018 125

126 24.08.2018 126

127 28.08.2018 127

128 06.09.18 128


129 12.09.18 129

130 12.09.18 130

131 12.09.18 131

132 14.09.18 132

133 15.09.18 133

134 15.09.18 134

135 17.09.18 135

136 19.09.18 136

137 23.09.18 137

138 23.09.18 138

139 23.09.18 139

140 23.09.18 140

141 24.09.18 141

142 01.10.18 142

143 05.10.2018 143

144 05.10.2018 144

145 11.10.2018 145

146 12.10.2018 146


147 13.10.18 147

148 13.10.18 148


149 14.10.18 149
150 15.10.18 150
151 15.10.18 151
152 15.10.18 152
153 18.10.18 153
154 19.10.18 154
155 19.10.18 155

156 20.10.18 156

157 21.10.18 157

158 20.10.18 158


159 21.10.18 159
160 22.10.18 160
161 23.10.18 161
162 24.10.18 162
163 24.10.18 163
164 24.10.18 164
165 25.10.18 165
166 25.10.18 166
167 27.10.18 167

168 28.10.18 168

169 28.10.18 169


170 28.10.18 170

171 29.10.18 171

172 29.10.18 172

173 29.10.18 173

174 30.10.18 174

175 30.10.18 175

176 30.10.18 176

177 02.11.18 177

178 02.11.18 178

179 02.11.18 179


180 02.11.18 180
181 03.11.18 181
182 09.11.18 182

183 10.11.18 183

184 10.11.18 184

185 12.11.18 185

186 14.11.18 186


187 15.11.18 187
188 17.11.18 188
189 22.11.2018 189
190 24.11.2018 190

191 24.11.2018 191

192 24.11.2018 192

193 24.11.2018 193

194 26.11.2018 194

195 27.11.18 195


196 30.11.18 196

197 01.12.18 197

198 02.12.18 198

199 05.12.2018 199

200 05.12.2018 200

201 05.12.2018 201

202 05.12.2018 202

203 05.12.2018 203

204 08.12.2018 204

205 08.12.2018 205

206 09.12.2018 206


207 10.12.2018 207

208 11.12.2018 208

209 12.12.2018 209

210 12.12.2018 210

211 14.12.2018 211

212 14.12.2018 212

213 14.12.2018 213

214 14.12.2018 214

215 14.12.2018 215


216 15.12.2018 216

217 16.12.2018 217

218 17.12.2018 218

219 17.12.2018 219

220 18.12.2018 220

221 18.12.2018 221


222 21.12.2018 222
223 22.12.2018 223
224 22.12.2018 224

225 27.12.2018 225

226 28.12.2018 226

227 01.01.19 227


228 01.01.19 228
229 01.01.19 229
230 04.01.19 230
231 06.01.19 231
232 08.01.19 232
233 09.01.19 233

234 10.01.19 234

235 12.01.2019 235

236 14.01.19 236

237 14.01.19 237


238 14.01.19 238
239 15.01.19 239
240 16.01.19 240

241 18.01.19 241


242 21.01.19 242
243 21.01.19 243

244 23.01.19 244

245 28.01.19 245

246 28.01.19 246

247 28.01.19 247


248 28.01.19 248
249 30.01.19 249
250 31.01.19 250
251 1.02.2019 251
252 1.02.2019 252
253 03.02.2019 253
254 03.02.19 254

255 07.02.19 255

256 08.02.19 256

257 09.02.2019 257

258 11.02.19 258


259 15.02.19 259

260 18.02.19 260

261 05.03.2019 261

262 08.03.19 262

263 09.03.19 263

264 12.03.2019 264

265 14.03.19 265

266 17.03.19 266

267 24.03.19 267

268 24.03.19 268

269 25.03.19 269

270 25.03.19 270

271 25.03.19 271

272 10.4.19 272

273 18.04.19 273

274 20.04.2019 274

275 21.04.2019. 275

276 21.04.2019. 276

277 21.04.2019 277

278 23.04.2019 278


279 24.04.2019 279

280 24.04.2019 280

281 24.04.2019 281

282 24.04.2019 282

283 24.04.2019 283

284 24.04.2019 284

285 25.04.2019 285

286 25.04.2019 286

287 25.04.2019 287

288 25.04.2019 288

289 09.05.19 289

290 10.05.19 290

291 10.05.2019 291

292 10.05.2019 292

293 10.05.2019 293

294 14.05.2019 294

295 15.05.2019 295

296 15.05.2019 296

297 20.05.2019 297

298 07.06.2019 298

299 07.06.2019 299


300 10.06.2019 300

301 15.06.2019 301

302 24.06.19 302

303 24.06.19 303

304 24.06.19 304

305 24.06.19 305

306 24.06.19 306

307 24.06.19 307

308 24.06.19 308

309 02.07.19 309

310 04.07.19 310

311 08.07.19 311

312 10.07.19 312

313 15.07.2019 313

314 15.07.2019 314

315 15.07.2019 315

316 15.07.2019 316

317 25.07.2019 317

318 29.07.2019 318

319 29.07.2019 319

320 31.07.2019 320

321 06.08.2019 321


322 07.08.2019 322

323 07.08.2019 323

324 26.08.2019 324

325 26.08.2019 325

326 29.08.2019 326

327 29.08.2019 327

328 29.08.2019 328

329 01.09.2019 329

330 11.09.2019 330

331 15.09.2019 331

332 19.09.2019 332

333 20.09.2019 333

334 22.09.2019 334

335 27.09.2019 335


Indian Oilt
(BOOT-3,
Paradip R

WORK REQUE

REQUEST DESCRIPTION DEPT

Leakage from Coolling Line of Fire engine 90-PA-103 Mechanical

FCV-3323 of MS-P CTMS hunting during vessel loading Instrumentation

Gland Leakage From Tower Monitor02 & 03 MOV Mechanical

MS(P) ctms steam-1 upstream HOV greasing Required Mechanical

Rusted Hose Box replacement required Mechanical

Drop wise leakage from MLA-028 QCDC during idle condition Mechanical

Tower monitor-1 & 4 MOV Gland Leakage Mechanical

Wind Shock to be Replaced of FWPH Mechanical

Tower Monitor Jockey Pump 101B NRV passing Mechanical

Hydrant Jockey Pump 102 A not started Electrical

DPK ctms Stream-4 Basket Filter chocked and to be repalced Mechanical

Leakage from high Vent point of Hydrant Network at Jetty Top Mechanical

No powwer Supply to HSD-E ctms Mov's Electrical

Leakage in HSD-D Prover upstream & downstream Flange Mechanical


Poor Illumination at VRU Area Electrical

Naphtha Stream -2 in CTMS minor leakage from Basket Filter overfill hole Mechanical

Tower Monitor-2 up/Down movement not working Electrical

HPU pump suction & discharge filter chocked, to be cleaned Mechanical

Tower Monitor Jockey Pump 101-A NRV passing Mechanical

Tower Monitor Jockey Pump 101A Emergancy push button not working Electrical

Tower Monitor-2 Left/Right movement not working Electrical

Tower Monitor-1 Jet/Fog movement not working Electrical

MOV-1302 at Battery Limit of Naphtha Line not functioning Instrumentation

Wind shock modification required at Substation Top Mechanical

leakage from MS(P) ctms stream-4 Basket Filter Cover overfill point Mechanical

Hydrolic leakage observed from MLA-04A QCDC hose Mechanical

Tower Monitor-2 Up/Down movement not working Electrical

Abnormal Sound during operation of TM network Jockey pump P-101B Electrical

MS-R Stream 2 not functioning properly Instrumentation

NRV passing of Jockey pump P-101B Mechanical

MLA-04B QCDC lock status not updating in DCS & LCP Instrumentation

Gangway Tower-1 Up/Down movement not working Electrical

SOJ back up process area Illumination not proper Electrical

HPU unit hydrolic oil quality found not proper,observed sluddgy & foreign
Mechanical
perticles in the oil.
HSD-E CTMS densitometer not working properly Instrumentation

Leakage observed from MLA-028 vaccum breaker during loading Mechanical

Tower Monitor-4 Left/Right, Up/Down movement not working Electrical

Gangway-2 not properly rested on supporting legs during perking position Mechanical

QRMH-4 locking plate of Hook-2 brocken due to high load devoloped


during vessel sail out time because of miscommunication in between Mechanical
mooring team & vessel crew.

All CTMS densitometer pump network 2" gate valve jamed Mechanical
DPK CTMS densitometer pump FT-210-3105 not giving proper reading at
Instrumentation
DCS.
Fire Bucket Stand modification Required Mechanical

Hydraulic leakage in gangway-2 during in & out movement Mechanical

Fire engine - 108 battery terminals found loose contact Electrical

Oil dripping in fire engine 103 fuel outlet during running Mechanical
Jockey Pump 101A/B for Tower Monitor is not starting from Loavl Control
Electrical
panel
Tower monitor -02 (JET) command is not working Electrical

Gangway-02 Telescopy in/out piston cover found damaged Mechanical

QRMH-05 capstan not working through padel push Electrical

Tower monitor-04 JET/FOG movement water flow is not working Mechanical

Tower monitor-01 Up /Down movement and Jet/ Fog movement is not


Mechanical
working

International shore connection HOV found jammed could not open Mechanical

During loading HSDD CTMS stream -3 flow indication was zero Instrumentation

In Gangway -1 hydrolic leakage found during left/right movements Mechanical

TM-4 left /Right movement not working Electrical


DPK CTMS Stream-2 to master stream MOV-3104 mechanical indication
Mechanical
not working
MLA-5A pre alarm hooter not working Instrumentation

locking plate of MD-1 hook-2 broken Mechanical

MLA-28 ERC ball valve passing and product accumulates in ERC down
Mechanical
stream

HSD-D fast loop pump is not working in remote and manual mode Electrical

MLA-28 vaccum breaker flange joint leakage when the MLA pressure
Mechanical
exceeds 2.0kg/cm2 during loading to MT Pacific Ziorocon

During De-berthing of MT PACIFIC ZIRCON Berthing dolphin 02 - hook01


Mechanical
locking plate broken

During trial of MLA 210-Z-04A freewheel alarm hooter sound is low Instrumentation

Naphtha CTMS PCV downstream HOV wheel is free revolving Mechanical

Naphtha CTMS PCV upstream HOV wheel is hard to open Mechanical

MLA-04B QCDC proximity switch failure indicator in LCP panel Instrumentation

MLA-04B hydraulic hose leakage found Mechanical

Level Gauge of Fuel Tank Fire Engine 106 & 107 is showing error Instrumentation

MLA 210-Z-04A Accumulatior pressure not holding Mechanical

Leakage from Storz connection Hov Mechanical


Ms-R CTMS ,Stream -4 Downstream,By-Pass and PCV downstream Hov
Mechanical
hard o open ,need greasing

Heavy corrision and unsafe condition occur in climb rod of Tower Monitor
Mechanical
ladder TM-01, TM-02, TM-03 & TM-04.

MLA 210-Z-028 accumuialtor lever bolt is broken Mechanical

Tubelight is not working at MCR east side exit gate Electrical

Pole light at HSD-D & MS-P CTMS is not working Electrical

NAPHTHA fast loop pump is not working in remote and manual mode Electrical

Leakage found from Naphtha 2nd stream busket filter cover Mechanical
Intercom no-102 Near Naptha CTMS 1&2 numbers not functioning on dail
Instrumentation
pad.
HSD-D Pig barrel pressure guage not going beyond 2.0kg/cm2 when
Instrumentation
pressurizing the pig barrel more than 5.0kg/cm2.
HSD-D pig barrel door, vent line 1/2inch valve not working. Mechanical

HSD-E Jetty top MOV-1008 selector switch for Auto/Manual not available. Instrumentation

HSD-E Jetty top MOV-1008 mechanically jammed . Mechanical

Propylene Liquid CTMS bypass HOV hand wheel is not available. Mechanical

Fire engine 104 cooling line Y strainer cap is opened Mechanical

MLA-04A 2nd satge alarm showing in DCS Instrumentation

MLA-05B hydraullic leakage from freewheel Solonied valve Mechanical

HSD-D PCV not taking commond from DCS(CTMS area) Instrumentation

Jetty Top Control Room Split AC outdoor unit fan is found broken Instrumentation

Base Monitor-04 HOV greasing required Mechanical

ESD-2 (2nd stage alarm) is showing in DCS of MLA-04A Instrumentation

MLAs alarm siren sound observed very low Instrumentation

Fire engine - 104 cooling line Y strainer cap is opened Mechanical

No illumunation(light failure) at QRMH-1 Electrical

Fire Engine -106 jerking observed during trail run in idle mode. Mechanical

No illumination (light problem) at FWPH staircase Electrical


MLA-04A Apex/slew movement alarm observed at LCP panel & 2nd
Instrumentation
satge alarm showing in DCS
Naphtha CTMS stream 04 FCV 3304 not opend fully after giving
Instrumentation
command 100%
Fire engine-103 RPM increasing Automatically in idle mode running time
& engine not stopped while press stop button and also emergency stop Mechanical
button.
MLA 04B Right & Left Hydraulic Hose Leakage observed during
Mechanical
Disconnection
Foam Pump 90-PA-202 are not discharging sufficient AFFF Mechanical
QRMH Electrical Hook Releasing operation not working
Electrical
QRMH02- Hook-1, QRMH 04 Hook-01 & QRMH 04 Hook-02
QRMH 05 Foot Capstan switch is not working Electrical

Foam Pump 90-PA-201 Frequent trip observed during Running time Electrical
MD-1 capston motor cover/canopy damage Mechanical

TM-4 left/right movement not working Electrical

NAOCL system shock dosing pump inlet & outlet line leakage Mechanical

MS-P fast loop pump is not running but flow is showing in FT (210-FT-
Instrumentation
3325)
HSD-D CTMS Stream-3 FCV (210-FV-3323) is not clossing fully Instrumentation
MLA-05A hydraulic oil line tube got punchered and hydraulic oil leakage
Mechanical
observeed
Proximity switch Failure of MLA 210-Z-028 Instrumentation

TM-2 left/right movement not working Electrical

BD-2, Hook-1 is not holding ropes properly Mechanical

TM-1, 2, 3 & 4 heavy rust found on ladder Mechanical

MLA-009 freewheel valve alarm status is showing 'Alarm' Instrumentation

Dropwise leakage from HSD CTMS Stream-03 DPT tubing point Instrumentation

Hydraulic oil leakage from MLA04A ERC ball valve hose Mechanical

Hydraulic hose found damaged for MLA 04A for lift, lower, forward &
Mechanical
backward movement

19OWS-DCS-1902 graphics in not updating properly Instrumentation

Unable to access to events of DCS from operator station Instrumentation

CBD pump-03 B discharge HOV is very tight Mechanical

MS-P battery limit analog pressure guage is not working Instrumentation

HSD-D ctms stream-3 not showing the flow in DCS Instrumentation

CBD level transmitter display reading not showing in field Instrumentation


TM-01 panel power supply isolated due to spark observed in RYB phase
Electrical
indicator
Jetty Top MS-P MOV'S 1011,1012 found product passing Mechanical

Canopy required for CBD panel & oil skimmer pump Mechanical

2" drain valve handle of all MLA riser to be replaced Mechanical

Proximity switch Failure alarm of MLA 210-Z-04A & 04B Instrumentation


jetty MCR DCS system hanged Instrumentation

HB-14 found door broken.need to maintenance Mechanical

AFMCP 406 & 407 FAULT ALARM OBSERVED. Instrumentation

Fire Engine 105 cooling line Y stainer leakage observed during trial run Mechanical
Naphtha CTMS stream -2 cummulative quantity not updatedin batch
Instrumentation
format
MS-R ctms mov-3604 not open in remote mode Instrumentation

HSD CTMS stream-3 PT-3023 reading is not showing properly Instrumentation

Gangway-1 hydraullic lock/release push button is very tight Electrical


Gangway -1 hydraulic line for manual operation leakage observed when
Mechanical
operated
Mechanical locking need of mla-009 not fully unlocked & sound observed
Mechanical
during left/Right movement

Painting rust preventation requirement of skimmer power pack system Mechanical


MS-R pig barrel pressure gauge not working & pressure constant at
Instrumentation
1.0kg/cm2
MS-R line MOV-1518 not working,unable to open Instrumentation

MLA-04B vacuum breaker NRV is not working Mechanical

BD-1, Hook-2 is not moving in right side. Mechanical

HSD-D STREAM-3 flowmeter not indicating flowrate. Instrumentation

Broken street light at backup area due to storm (Cyclone Titli) Electrical

Jockey Pump 101A/B & 102A amps meters is not working. Electrical
MLA-09,4A & 5A raiser bonding wire corroded and broken. Electrical

Hose reel No-06 drum has been damaged, need to be replaced. Mechanical
Leakage observed from Naphtha CTMS stream-2 Basket Filter top cover Mechanical
Foam supply pump 90-PA-201 main power cable laying outside Electrical
Hose reel box in front of VRU is damaged Mechanical
All MLA of Jetty top area Earthing/bonding cable heavy rusting/ corroding
Electrical
done
MLA-04B vaccum breaker rope broken while arresting the minor dripping
Mechanical
of product
Water dripping from split AC at FWPH panel room Electrical
Tower Monitor up/down movement not working Electrical

TM-04 mov-6 not working in remote mode Instrumentation

Hydralic lock/release push button journal. Also Emergency release knob


Electrical
missing
TM-02 water MOV-4 is not working both local & remote mode Instrumentation
TM-02 Left movement is not working Electrical
MLA-04A MS-P MOV-1012 torque fault ocurred Instrumentation
90-PA-105 RPM indication not working properly Electrical
Hydraulic leakage observed from mla-04B Mechanical
Jumbo curtain mov-07 gland leakage Mechanical
MLA-05A Inboard arm mechanical lock jam Mechanical
MLA-04A ERC ball valve hydraulic hose leakage Mechanical
Jetty Top Control Room PMS not working Instrumentation
MLA 210-Z04B Lift Lower Cylinder PRV Damaged and Hydraulic oil
Mechanical
Leakage observed
Proximity switch failure in MLA27-showing apex/slew alarm in arm lock
Instrumentation
condition. MLA4B & MLA5B - QCDC clamped
Jet/Fog movement of TM-4 Mechanical
Jet/Fog water flow not properly due to stuck of foraign material (BM-1) Mechanical
Mechanical /
TM-2 Up/Down movement not occurs
Electrical

Mechanical /
TM-1 Left/Right movement not occurs
Electrical
Naphtha Stream-4 MOV-3309 feedback problem Instrumentation

Fire engine 108 jerking observed during trial run Mechanical

DPK CTMS Stream -03 Leakage observed from Basket filter top cover Mechanical

Fire engine 106 over speed Trip occurred , engine speed 842 rp, Mechanical
WFM 04 Isolated valve, nozzle found tightened. WFM 12 eft Right
Mechanical
movement found tightened
Double hydrant near TM-3 valve tightened, need to be greasing Mechanical
All CTMS HOV's gear box empty & no grease observed in vapour
Mechanical
eliminator, PCV up/downstream HOV's
Street light way to FWPH roadside street light is not working Electrical
Naphtha CTMS, Master stream MOV-3302 is greasing requirement Mechanical
Naphtha PCV loop u/s HOV tight greasing required Mechanical
Propylene Vapour line HOVs servicing required at battery limit area as
Mechanical
the HOVs are very hard to operatre .
Propylene liquid and vapour lines HOVs handles to be fitted at battery
Mechanical
limit area
Tower Monitor-01,02 & 04 MOV's not working Instrumentation
Tower Monitor Network furthest point PG is not working at Jetty Berth
Instrumentation
area
MLA-4A hydraulic oil hose at ferrule got punctured Mechanical
Naptha CTMS Stream-4 MOV(210-MOV-3309) not opeerating from DCS Instrumentation
Leakage observed from FE-104 &106 sea cooling water line during run
Mechanical
operation
FE-107 oil leakage from lube oil filter Mechanical

FE-105 Y-strainer & FE-104 sea water cooling line leakage Mechanical

MLA 210-Z-028 lock/unlock sensor is not working properly Instrumentation


HSD-D CTMS PCV(PV-3025) not opening 100%. When open command
Instrumentation
given to 100%, it is opening only 94%
MLA 04B vaccum breaker dropwise leakage observed during loading of
Mechanical
MT HARMONY
MLA-04B QCDC feedback not coming in DCS Instrumentation
Fire engine 105 battery got discharged and fire engine not started Electrical

Pathway lighting from QRMH-06 to QRMH-05 not glowing properly. Electrical

Less illumination at MCR (visual observation) Electrical

FWPH diesel supply pump not working properly Mechanical


DPK CTMS stream-03 MOV no-3106 actuator is not coupled with valve
Mechanical
stem
DPK PIG Receiver isolation valve MOV-1298 selection switch not working Instrumentation
NAPTHA PIG Receiver isolation valve MOV-1302 selector switch not
Instrumentation
working

Low illumination at Foam tanks Electrical

Dropwise leakage from HSD CTMS By Pass Hov's (Duty meter- 1 ,2 & 3) Mechanical

Hydraulic Hose Leakage from Propylene Vapour MLA-027 Mechanical

HSD-D battery limit MOV-1549 is not closing in local/remote mode only


Instrumentation
6% closing
Emergency exit light at MCR is not working Electrical
90-PA-105 Fire engine diesel tank level instrument not showing exact
Instrumentation
value.
MLA-04A QCDC , found abnormal sound during movement. Mechanical

Jetty top MOV-1012 not opening/closing compeletly through actuator


Mechanical
found .

Foam Feed Local control Panel & Remote Control panel power on
Electrical
indication is not working
Foam Feed Local control Panel & Remote Control panel Ammetre is not
Electrical
showing current

90-PA-106 Pump Gear Oil Level is low Mechanical

90-PA-106 Pump Engine oil quantity to be checked and filled Mechanical

Both of the 2''-Nitrogen line Gate Valve passing in MLA-04B line Mechanical
Diesel top up pump not working at FWPH Mechanical

Coupling guards to be fitted for condensor pumps in chiller room Mechanical

QRMH-2 &4 hook no-01 mechanically jammed during release operation Mechanical

All PCVs Tuning to be done Instrumentation

Alarms configuration in DCS system Instrumentation

TM-04 water MOV-06 feedbak error in DCS Instrumentation


FE-105 diesel tank LT showing abnormal valve Instrumentation
Emergency fire siren
DPK CTMS-stream-04 chocked during loading. Mechanical

Fire engine 104 Diesel tank sweating and dropwise leak observed Mechanical
Leakage observe union coupling of FE-106 during FE running in idle
Mechanical
mode
Gangway 02 Luffing Up/Down hydraulic hose rapture Mechanical
Gangway 02 Telescopic in/out Switch found not working Electrical
CBD P-03B Discharge HOV Handle being Free , Fixing of Handle is
Mechanical
required
During Trial run of Gangway-01,hydraulic hose reptured Mechanical
During HSD loading Stream-3 FT-3023 no flow shows Instrumentation
Gangway-2 Bulwalk up/down hydraulic hose reptured Mechanical
Gangway-1 &2 Bulwalk Cylinder Piston protected guard need to change Mechanical
FE-105 batery terminal repairing required as during trial run sparking
Electrical
observe due at the time of FE-105
MS-P Jetty top MOV(1012) Could not start locally before vessal loading
Instrumentation
(torque trip alarm)
Fire Engine-108 not started , Low Lube Oil Pressure showing in panel Mechanical

Sample room lighting requirement Electrical


Battery room Lighting at gate is not properly glowing Electrical
Tower Monitor-01 Up & Down movement not working from LCP Electrical
CBD Pump 003B Discharge HOV Hand Wheel Check Nut grab screw is
Mechanical
missing
MD-01 Hook-02 release hook is not working Electrical
ATF CTMS Stream 03 MOV 3103 actuator flange leakage observed Mechanical
ATF CTMS Stream 03 down stream HOV gland leakage observed Mechanical

ATF stream-02 FT not showing flow during loading Instrumentation


HSD-E PCV downstream HOV bonding wise one end is observed in open
Electrical
condition
HSD-E PCV downstream HOV gland dropwise oil leakage observed Mechanical
HSD CTMS Stream-2 & 3 bypass HOV dropwise leakage observed from
Mechanical
gland
Slight dropwise leakage observed from HSD-D CTMS D/S HPV Mechanical
FT-3323 stream-3 of ms-p(CTMS) ,No flow value shown during Vessel
Instrumentation
Loading
CBD Pump-03B Discharge hov handwheel free Mechanical
MOV-1013 (J/T) thermostate trip found during open. Instrumentation
5 no. of street light found off condition infront of MCR. Electrical
Pole light near Propylene Vapour PIG launcher is not working Electrical
MS-P line CTMS Stram 02 & Stram 03 DPT Top off point flange leakage ob Mechanical
MS-P line CTMS Stream 03 up/downstream DPT Top off point flange
Instrumentation
leakage observed during Loading
Naphtha stream -02 MOV-3305 not opening in Remote and local beyond
Instrumentation
19%

Leakage observed from HSD-E stream-04 Basket filter top surface area Mechanical

MLA-027 & 026 nitrogen leakage from both of the QCDC flange Mechanical
MLA-009 ERC ball valve proximity sensor failure. Instrumentation

Drop wise leakage observed from 04A mla vent ppoint during loading
Mechanical
(when pr. Reached more than 1.65kg/cm2)

Fire engine 104 dropwise leakage in diesel return line at near tank during
Mechanical
trial run
In between, HSD-D ctms d/s and pcv u/s HPV minor leakage observed
Mechanical
during to MT JAG AANCHAL
Jetty top MOV-1007 (HSD-D) is physically fully closed but in actuator
Instrumentation
showing 18% open, Limit setting to be done.

Earthing cable found disconnected in both motors used to fill diesel tanks
Electrical
at FWPH

Tower Monitor-1 up movement is not working Electrical


Fire Engine 105 sea water cooling water return line leakage observed
Mechanical
during trial run
DPK CTMS stram-2 FT-3102 is not showing flowrate during ATF loading
Instrumentation
to MT DAWN HARIDWAR
MLA-04B Hydraulic hose puncher Mechanical

FWPH Building UPS Room Ac's not working Electrical

DPK CTMS Stream-03 mov 210-Mov-3307 product dripping observed mechanical

Actuator lube oil dripping observed from DPK CTMS Stream-02 Mov &
Instrumentation
Stream-04 MOV

TM-2 & TM-4 up-down movement not working Electrical

No illumination (light failure at back up area Electrical

QRMH-01 Hool-01 Locking clit found damaged need to be replace Mechanical


MS-P CTMS Stream-03 TT (9210-TT-3325) gives over range alarm at
Instrumentation
29.5deg.c

During loading MLA-04B ERC ball valve alarm coming in DCS Instrumentation

DPK CTMS stream-3 MOV's acctuator running but MOV not operating mechanical

DPK CTMS sream-02 FT(3102) is not showing flowrate during MT HARSI


Instrumentation
SAGAR loading
DPT-3602 of MS-R CTMS stream-2 in high alarm as reaches to
Instrumentation
0.75kg/cm2

During bullwark ladder down also luffing down in gangway-2 Electrical

Gangway-2 bulwark hydraulic hoses are scattered mechanical

During MLA-05A trial run, return hydraulic line rupture near abinet tubing
mechanical
line connection

Propylene liquid CTMS stream-2 basket filter gets chock mechanical

Liquid propylene CTMS-3 basket filter chock mechanical

MOV-3029 HSD stream-4 oil dripping observed from actuator Instrumentation

HSD stream-4 PT ferrule joint minor oil dripping observed Instrumentation

HSD-D isolation valve-2 MOV-1549 freeness of wheel is tight (greasing


Instrumentation
required)

DPK CTMS stream-4 FT not showing flowrate during loading to MT


Instrumentation
TORM HELVIG

M/A-04B nitrogen valve damage Mechanical

MLA-04B QCDC un flanged hydraulic hose damage Mechanical

Liquid Propylene CTMS, Stream-2 ,MOV-4305 not operated in Local as


Instrumentation
wellas in Remote mode

FCV-4301, neither open or close during operation Instrumentation

During Gangway-2 put on Vessel Deck,Luffing IN Button is not working Electrical

During Safety Walk Through, it was found that fire bucket painting got
Mechanical
damaged

From DCS LT's reading is not matching to Fuel tank's manual reading Instrumentation

From hydraulic cyliner of Luffing Up & down of Gangway Tower 02 minor


Mechanical
oil sweating observed
During MLA Dis-connection, suddenly Hydraulic hose of MLA 04A QCDC
Mechanical
raptured

Jockey pump 101A not started in Auto mode Electrical

Jockey pump 101B leakage observed from discharge PG point joint Mechanical
MLA-04B QCDC feedback not coming in DCS Instrumentation

MLA-009 minor sweating observed at vacuum breaker Mechanical

hydraulic hose rectification & top up of hydraulic oil in gangway-1 Mechanical

Fire hydrant line leak arresting of equipments Mechanical

Emergency siren of both gangway-1 & 2 not working also fixing required
Electrical
for both

PA-1007(At jetty top) mike sound not clear. Electrical

Field telephone-106 not working. Electrical

Field telephone -101 gland & wire remove from system. Electrical

Beacon siren of MLA'S at jetty top missing. Instrumentation

MLA-04B Vacumm breaker leakage observed Mechanical

MD-03, HOOK-02 FOUND JAMMED Mechanical

MLA-04A QCDC & ERC of hydraulic hose damaged at outboard arm end. Mechanical

Post / Pole lights S.No. 115,116,112,124 and 126 found not wokring on
Electrical
MD-04 platform area.

FE 106 not stopped by its local panel Mechanical

FE 106 not stopped by its enginel panel Mechanical

Leakage found from hydroulic line of gangway 2,bulkwalk ladder cylinder Mechanical

Gangway 1 bulkwalk ladder switch to be provided at gangway platform


Electrical
panel and lock/release switch to be repair at bulkwalk ladder side

FT-3323 stream-3 of MS-P (CTMS) ,High flow value fluctuation shown


Instrumention
during Vessel Loading

FE-105 Water Coollent Line Punture Mechanical

FE-107 not stopping from local panel Mechanical

Hydraulic Oil leak observed from out board arm of MLA-09. Mechanical

MLA-04A QCDC and backward hydraulic hose ruptured during


Mechanical
connection.
MLA 04A hydraulic oil dripping observed at spool piece solenoid valve. Mechanical

QRMH-5 capston not working. Electrical

QRMH BD-1 & BD-2 BOTH HOOK-01 not working electrically Electrical

QRMH BD-1 & BD-2 BOTH HOOK-01 Mechanically jammed Mechanical

MLA-04A 2" draining line first isolation HOV from mla side leakage
Mechanical
observed from gland.

MLA-05A,009,028,026,027,01A drain valve handle not available mechanical

MLA-05A,009,028,026,027,01A drain valve handle corroted mechanical

BD-2 HOOK-2 not realese from panel Electrical

Fire engine-103 fuel tank gauge flange sweating observed Mech & Inst

FE-105 Water Coollent Line Leakage found Mechanical

MLA 04A QCDC poximitysensor not working instrument

MLA04B XZV 026 not closing INSTRUMENT

MLA-028 minor sweating observed from vacum breaker Mechanical

GANGWAY-2 selector switch (luffing up/down ) mallfunctioning occure Electrical


ndian Oiltanking Limited
(BOOT-3, South Oil Jetty)
Paradip Refinery, Paradip

WORK REQUESTS ( From-April 2018)

SOUTH OIL JETTY

ACKNOWLEDGED
OBSERVED BY ACTION TAKEN
BY
Leakage arrested from Fire Engine 90-PA-103 cooling line by
Alton WR Praveer Sinha
doing hot welding job

Pallab M Abhishek Sati Rectified by OEM vendor

TM-02 & 03 main water supply MOV gland leakage arrested


B Raju Praveer Sinha
by tightening gland packing

Rahul Ray Praveer Sinha MS-P CTMS Stream-1 up stream HOV made free by greasing

Ramesh J Praveer Sinha All rusted hose boxes replaced by new fiber made hose boxes

MLA-028 QCDC "O" ring replaced with a new one & coupled
Pranjal B Praveer Sinha
jaws are adjusted, No leakage observed
TM-01 & 04 main water supply MOV gland leakage arrested
Pranjal B Praveer Sinha
by tightening gland packing

Manoj D Praveer Sinha FWPH wind shock replaced with new one

TM Jockey pump 101B NRV problem attended &


Pranjal B Praveer Sinha
troubleshooting done, made ok

Raushan Kumar Pratap Patra Troubleshooting done & made ok

DPK CTMS Stream-4 Basket filter open, filter element cleaned


D K Saha Praveer Sinha
properly for foreign perticles and refitted, made ok
Leakage arrested by placing a new vent cap with a new vent
Rahul yadav Praveer Sinha
valve at Hydrent Network at Jetty Top area

Harshit M Pratap Patra control wiring changed

Leakage arrested by tightening HSD-D both Up & Down


Santhosh Reddy Praveer Sinha
stream flage nut bolts.
Pallab M Pratap Patra

B Raju Praveer Sinha Leakage arrested by tightening basket filter top cover bolts .

Rahul yadav Pratap Patra Reset push button

Cleaned both chocked Suction & discharged filter of HPU by


Ramesh J Praveer Sinha
instrument air flushing.
TM pump P-101A attended for NRV passing & NRV found
Raushan Kumar Praveer Sinha
damage, consulted with IOCL for replacement of newer one.

Pallab M Pratap Patra changed with new push button

Ramesh J Pratap Patra reset contactor

Raushan Kumar Pratap Patra Rust clearing

Alton WR Abhishek Sati Rectified during PM

Manoj D Praveer Sinha Damage Wind shock replaced with a new one at sub station.

Leakage arrested by tightening basket filter top cover all bolts


D K Saha Praveer Sinha
& replacing 2nos of damage bolts with new one .

Raushan Kumar Praveer Sinha MLA-04A QCDC damage hose replaced with new one.

Pallab M Pratap Patra Troubleshooting done

B Raju Pratap Patra Troubleshooting done

Pranjal B Abhishek Sati Rectified after maintenance

TM Jockey pump 101B NRV problem attended &


Rahul Ray Praveer Sinha
troubleshooting done, made ok

Rahul yadav Abhishek Sati Rectigfed from field

Pranjal B Pratap Patra Rust clearing

Rahul Ray Pratap Patra

All old detoriated quality hydrolic oil drain out & topped up with
Pranjal B Praveer Sinha
new good quality oil in HPU cabinet.
B Raju Abhishek Sati To be commisioned by OEM

Leakage arrested by MLA-028 vacum breaker valve inspection


Alton Wr Praveer Sinha
& servicing and same is re installed

B Raju G P Sahoo Troubleshooting done

Pranjal B

Brocken plate is replaced with new one & plate properly


Raushan Kumar Praveer Sinha
installed at Hook-2 of QRMH-4

Pranjal B Praveer Sinha

Pranjal B Abhishek Sati spare to be uesd for replacement after finalization with IOCL

Manoj D Praveer Sinha Fire bucket stand modified properly placed.

B Raju Praveer Sinha End push washer & o-Ring replaced and leakage arrested

Ramesh J G P Sahoo Troubleshooting done

Alton WR Praveer Sinha oil dripping arrested

Raushan Kumar Pratap Patra Troubleshooting done

Manoj D Pratap Patra Troubleshooting done

Manoj D

Harshit M G P Sahoo Troubleshooting done

Pranjal B Praveer Sinha Foreign material removed from jet/fog nozzele

Rahul yadav Praveer Sinha All unwanted material removed from TM-04 jet fog movement

Jet fog gearbox found bend, make shaft & support guide
Pranjal B Praveer Sinha
straight by help of lathe machine.

Alton WR Abhishek Sati Troubleshooting done & flow incation now showing

Sanjib S Praveer Sinha Troubleshooting done & leakage arrested

Avdhesh singh G P Sahoo Troubleshooting done

Avdhesh singh Praveer Sinha Troubleshooting done & made ok


B Raju Abhishek Sati checked and found in working condition

Alton WR Praveer Sinha Broken locking plate of MD-1 hook-2 replaced with new one

Rahul Ray

Rahul Ray G P Sahoo

B Raju Praveer Sinha Troubleshooting done & made ok

Rahul Ray P Sarvanan Locking plate replaced

Rahul Ray Abhishek Sati Checked & found ok

Avdhesh singh

Avdhesh singh Praveer Sinha Greesing done and HOV is free now

Ramesh J Abhishek Sati checked and found ok

Hydraullic hose replaced fpor MLA-04B left right function found


Rahul yadav Praveer Sinha
ok

Pallab M Abhishek Sati checked and found in working condition

Pallab M

Raushan Kumar

Raushan Kumar

G.P.Sahoo

Pallab M Praveer Sinha Handle fixed and function can be done with the fixed handle

Pallab M Pratap Patra Tube light changed

Pallab M Pratap Patra

Rahul Ray Pratap Patra

Pallab M

Santhosh Reddy Abhishek Sati spare scope with iocl


Santhosh Reddy Abhishek Sati Troubleshooting done & made ok

Santhosh Reddy

Ramesh J Abhishek Sati spare scope with iocl

Ramesh J

Rahul Ray

Pallab M Praveer Sinha stariner cap fixed

Rahul yadav Abhishek Sati checked and made ok after sensor cleaning

Pallab M Praveer Sinha Change the repturehose by new one

Ramesh J Abhishek Sati Rectified from control room

Rahul yadav

Sanjib S Praveer Sinha Greasing done

Rahul yadav Abhishek Sati Rectified from Field

Santhosh Reddy Abhishek Sati Spares required & spares are under IOCL scope

Rahul Ray Praveer Sinha cap closed at the end flange

Narasinga Rao B Pratap Patra Troubleshooting done

Rahul Ray

Pallab M Pratap Patra Trouble shooting done

Pallab M Abhishek Sati

Ramesh J Abhishek Sati checked from field valve is mechanically stuck

Problem attend in presence of vendor, vendor suggested


Ramesh J Praveer Sinha
without placing PWM card don’t operate FE-103

Narasinga Rao B Praveer Sinha Hose replaced with new one & leak arrested

Pallab M

Avdhesh singh Pratap Patra Cleaned with CRC, now its ok

Avdhesh singh Pratap Patra Cleaned with CRC, now its ok

Pallab M Pratap Patra Meggering done. Now ok


Pranjal B Praveer Sinha New motor canopy fabricated & placed properly

Pranjal B Pratap Patra Cleaned with CRC, now its ok

Pranjal B Praveer Sinha Leakage arrested for shock dosing inlet & outlet line

Pallab M Abhishek Sati Spare parts not available, Follow up under progress with iocl

B Raju Abhishek Sati calibrated & found ok

Santhosh Reddy Praveer Sinha Damaged tube replaced with new one

Avdhesh singh Abhishek Sati Rectified from field

Narasinga Rao B Pratap Patra Cleaned with CRC, now its ok

B Raju

Pratap P

B Raju Abhishek Sati Rectified from field

Pranjal B Abhishek Sati Rectified from field

Ramesh J Praveer Sinha Damage hydraulic hose replaced with new hose

Raushan Kumar Praveer Sinha Damage hydraulic hose replaced with new hose

Issue mailed to IOCL for Check by OEM. follow up under


Rahul yadav Abhishek Sati
progress
Raushan Kumar Abhishek Sati Rectified by Honeywell

Avdhesh singh

B Raju Abhishek Sati checked and found ok

B Raju Abhishek Sati Rectified and checkedusing simulation

Alton WR Abhishek Sati checked and found ok

Ramesh J Pratap Patra LED ( R,Y,B) changed with new

B Raju

Raushan Kumar Praveer Sinha checked and found ok

Raushan Kumar

Rahul yadav Abhishek Sati MLA-04B ok & MLA-04A ESD-2 sensor fault rectified
informed to IOCL-fixed by honeywell rectified loose cabling &
Pranjal B Abhishek Sati
checked by simulation

Manoj D Praveer Sinha

Alton WR Abhishek Sati Checked & rectified

Ramesh J

Sanjib S Abhishek Sati Checked & tested by creating new batch and found ok

Sanjib S Abhishek Sati Checked Local & Remote mode operation and found ok

Raushan Kumar Abhishek Sati Zero calibration performed & now showing working properly

Avdhesh singh Pratap Patra

Raushan Kumar

Raushan Kumar Praveer Sinha complinced

Raushan Kumar Praveer Sinha Painting done

Santhosh Reddy Abhishek Sati PG maintenance performed. Now working properly.

B Raju Abhishek Sati Rectified from Field

Pallab M Praveer Sinha complinced

Pallab M Praveer Sinha Rectified

B Raju Abhishek Sati Rectified from field Checked & found ok

Ramesh J Pratap Patra Dismantle the broken light

Santhosh Reddy Pratap Patra


Santhosh Reddy Pratap Patra Replaced with new one

Manoj D Praveer Sinha Replaced hose reel


Alton WR
Pranjal B Pratap Patra Cable joint & termination done
Pranjal B Praveer Sinha New drum installed for hose box
Pranjal B Pratap Patra Earthing cable replaced with new
Santhosh Reddy Praveer Sinha New D-shackle replaced
Pallab M Pratap Patra Cleared the water choked area
Raushan Kumar Pratap Patra Cleaned with CRC, now its ok
Checked from field. Remote command not working from LCP,
Santhosh Reddy Abhishek Sati
to be checked by Elect Dept.

Raushan Kumar Pratap Patra Rectified

Pallab M Abhishek Sati Checked & rectified


Pallab M G P Sahoo Cleaned with CRC, now its ok
Pallab M Abhishek Sati Valve found getting struck during operation
Avdhesh singh Pratap Patra
Avdhesh singh Praveer Sinha Leak arrested
Pallab M Praveer Sinha Leak arrested
Pallab M Praveer Sinha Rectified
B Raju Praveer Sinha Rectified
Rahul Ray Abhishek Sati Communicated with IOCL
Santhosh Reddy Praveer Sinha New PRV Installed

B Raju Abhishek Sati Rectified from field

Praveer Sinha Jet/Fog movement for TM-4 troubleshooting done


Raushan Kumar Praveer Sinha Foreign particles removed

Pranjal B G P Sahoo Cleaned with CRC, now its ok

Raushan Kumar G P Sahoo Cleaned with CRC, now its ok

Alton WR Abhishek Sati Rectified from Field

Rahul Ray Praveer Sinha Rectified in the presence of vendor

Rahul Ray Praveer Sinha Rectified checked & found ok

Pallab M Praveer Sinha Rectified & fueling found satisfactory

Rahul yadav Praveer Sinha

Manoj D Praveer Sinha Rectified in the field

Raushan Kumar Praveer Sinha Greasing done


Alton WR Pratap Patra Clean termination of JB
Sanjib S Praveer Sinha Greasing done
Ramesh J Praveer Sinha Greasing done

Santhosh Reddy Praveer Sinha Valves servicing done

Santhosh Reddy Praveer Sinha Under scope of IOCL

Alton WR Abhishek Sati rectified all faults

Pallab M Abhishek Sati Isolation valve found stuck


B Raju Praveer Sinha Damaged hose replaced with new one
Santhosh Reddy Abhishek Sati rectified from control room panel
Rahul Yadav Praveer Sinha Leakage arrested using welding & M seal
Ramesh J Praveer Sinha M seal applied and leak arrested

Ramesh J Praveer Sinha Leakage arrested using welding & M seal

Santhosh Reddy Abhishek Sati rectified from field

Santhosh Reddy Abhishek Sati PCV re-calibrated, checked after calibration & found ok

Rahul Ray Praveer Sinha Re-installed After Cleaning and Servicing

Ramesh J Abhishek Sati rectified from field


Santhosh Reddy Pratap Patra Boost chrged,now ok

Raushan Kumar Pratap Patra Troubleshooting of Lighting done.

Pallab Manna Pratap Patra 06 nos tubes changed

Pranjal B Praveer Sinha New discharge valve installed & now it is working properly

Avdhesh singh

Avdhesh singh Abhishek Sati Lubricated & freed the selector switch

Avdhesh singh Abhishek Sati Lubricated & freed the selector switch

08no's tube lights fixed out of 12 no's, Remaining will be fixed


Ramesh J Pratap Patra
later, no access to fix now.

Raushan Kumar Praveer Sinha Leakage arresting by tighten all the Gland

Raushan Kumar Praveer Sinha Replaced raptured hose

Ramesh J Abhishek Sati Checked & found ok


Gautam Sil Pratap Patra New 6 no's spare fixture used, ok now

B Raju Abhishek sati Rectified from field

B Raju Praveer Sinha Proper venting done for QCDC cylinder and grease applied

B Raju

Pallav M Pratap Patra changed with new LED

Pallav M Pratap Patra

Sanjib S

Sanjib S

Pranjal B Praveer Sinha Replace both of the Gate Valve by newly Installed
Raushan Kumar Praveer Sinha

Santhosh Reddy Praveer Sinha New coupling guard fixed for both condenser pumps

Rahul Ray Praveer Sinha Lubrication and oiling done

Narasinga Rao B Abhishek Sati Tuning of all PCV is completed

Narasinga Rao B Abhishek Sati Completed for commissioned alarms

Narasinga Rao B Abhishek Sati Rectified from field


Pranjal B Abhishek Sati Rectified from field
Pranjal B
Ramesh J Praveer Sinha Basket filter opening, cleaning & boxup done

B Raju

Rahul yadav

Rahul Ray P Sarvanan Hose Changed


B Raju Pratap Patra Repaired Nob
Sanjib S Praveer Sinha Maintenance of Handle completed
Rahul yadav Praveer Sinha New Hydraulic Hose replaced
Raushan K Abhishek Sati
Rahul Ray Praveer Sinha Replaced with new one
Narasinga Rao B Praveer Sinha

Sanjib S Pratap Patra New earthing terminal fitted

Rahul yadav

Sanjib S

Raushan K Pratap Patra


Raushan K Pratap Patra
Narasinga Rao B Pratap Patra Cleaned with CRC, now its ok
Pallab M

Rahul Y
Narasinga rao B Praveer Sinha Leakage arresting by tighten all the flang bolts
Narasinga rao B Praveer Sinha Leakage arresting by tighten all the Gland

Rahul Y Abhishek Sati Checked & rectified card issue from field

Narasinga Rao B Pratap Patra replaced with newnut bolt

Pranjal B Praveer Sinha Leakage Arrested

Pranjal B Praveer Sinha Leakage Arrested


Pranjal B Praveer Sinha Leakage Arrested
Raiushan Abhishek Sati Checked & rectified card issue from field
Raushan

Rahul Y
Rahul Y
Pallab M
Avdhesh singh

Avdhesh singh

Raushan

Santhosh Reddy Praveer Sinha Leakage arresting by tighten all the nut bolts

Raushan K
Rahul Ray

Rahul Y Praveer Sinha Leakage arresting done by Mechanical team

Rahul Ray

Rahul ray

Raushan K

Pranjal B

Rahul Y

Santhosh R

Rahul Ray

Sanjib Show

Santhosh reddy

Santhosh reddy

Santhosh reddy

Avdesh singh

Narasinga Rao B

Santhosh Reddy Praveer Sinha Locking clit replaced

Raushan K

Ramesh J

sanjib show

sanjib show
Raushan K

Raushan K

Raushan K

Raushan K

Pranjal B

Pranjal B

Ramesh J

Ramesh J

Ramesh J

FT instrumentation card loosen from its position due to


Rahul Ray SS Malla
vibration, card was placed on its position & tightened properly

Ramesh J Praveer Sinha

Ramesh J Praveer Sinha Damaged hose replaced with new one


Raushan
SS Malla After Freenesss of Gear,MOV-4305 Working properly
Kumar
Raushan
SS Malla Calibration done by Instrument department
Kumar
Raushan
Pratap Patra Replaced the swithch by new one
Kumar

Rahul Yadav Praveer Sinha periodically painting to be done

Rahul Yadav SS Malla Calibration done by Instrument department

Rahul Yadav Praveer Sinha Preventive maintenance done by Mechanical department

Raushan
Praveer Sinha (hydraulic hose) relaced with new one
Kumar
Ramesh J

Ramesh J
Ramesh J SS Malla Rectification done

Ramesh J

Raushan K Gautam Sil replaced with new & hydraulic oil top up done

Raushan K Gautam Sil

Raushan K Pratap Patra

Raushan K Pratap Patra

Gautam Sil Pratap Patra

Gautam Sil Pratap Patra

Gautam Sil Pratap Patra

Avdhesh singh

Avdhesh singh Biplab mandal

Ramesh J Bharath R Damaged hose replaced with new one

Bharath

Sanjib S

sanjib s

sanjib s

Darshan P

FT instrumentation card loosen from its position due to


K Rana Ratnakar
vibration, card was placed on its position & tightened properly

Sanjib S

Sanjib S

Bharath

Darshan P Praveer Sinha Damaged hose replaced with new one


Ramesh J

Santosh R Pratap Patra

Avdhesh singh Pratap Patra Problem rectified

Avdhesh singh Praveer Sinha Both hooks made free

Avdhesh singh

Ramesh J

Ramesh J

K. Rana

Narasinga Rao B Ratnakar & Biplab Damaged Gas cut replaced with new one

Rahul M

Ramesh J Ramesh J Replace with new one

Ramesh J

Ramesh J

Rahul m
CLOSED By Checked By
ACTION TAKEN
Name Date Name
ed from Fire Engine 90-PA-103 cooling line by
Praveer Sinha 02.04.18 Rahul Y
ng job

EM vendor Abhishek Sati 06.04.18

ain water supply MOV gland leakage arrested


Praveer Sinha 09.04.18 Raushan Kumar
land packing

tream-1 up stream HOV made free by greasing Praveer Sinha 07.04.18 Santhosh Reddy

boxes replaced by new fiber made hose boxes. Praveer Sinha 07.04.18 Santhosh Reddy

C "O" ring replaced with a new one & coupled


Praveer Sinha 18.05.18 Pranjal B
ed, No leakage observed
ain water supply MOV gland leakage arrested
Praveer Sinha 10.04.18 Santhosh Reddy
land packing

ock replaced with new one Praveer Sinha 10.04.18 Santhosh Reddy

mp 101B NRV problem attended &


Praveer Sinha 12.05.18 Pranjal B
g done, made ok

g done & made ok Pratap Patra 10.04.18 Rahul Ray

eam-4 Basket filter open, filter element cleaned


Praveer Sinha 12.04.18 Pranjal B
eign perticles and refitted, made ok
ed by placing a new vent cap with a new vent
Praveer Sinha 13.04.18 Raushan kumar
nt Network at Jetty Top area

hanged Pratap Patra 15.04.18 Pallab M

ed by tightening HSD-D both Up & Down


Praveer Sinha 16.04.18 Pallab M
ut bolts.
ed by tightening basket filter top cover bolts . Praveer Sinha 25.04.18 Harshit Mishra

tton Pratap Patra 16.04.18 Pranjal B

chocked Suction & discharged filter of HPU by


Praveer Sinha 18.04.18 Ramesh J
lushing.
01A attended for NRV passing & NRV found
Praveer Sinha 11.05.18
ulted with IOCL for replacement of newer one.

new push button Pratap Patra 22.04.18 Pallab M

Pratap Patra 25.04.18 Rahul Y

Pratap Patra 04.05.18 Raushan Kumar

g PM Abhishek Sati 10.05.2018

shock replaced with a new one at sub station. Praveer Sinha 07.05.18 Manoj D

ed by tightening basket filter top cover all bolts


Praveer Sinha 08.05.18 B Raju
os of damage bolts with new one .

C damage hose replaced with new one. Praveer Sinha 10.05.18 Raushan Kumar

g done Pratap Patra 11.05.18 Pranjal B

g done Pratap Patra 13.05.18 B Raju

maintenance Abhishek Sati 20.05.2018

mp 101B NRV problem attended &


Praveer Sinha 12.05.18 Rahul Ray
g done, made ok

field Abhishek Sati 20.05.2018

Pratap Patra 18.05.18 Pallab M

ed quality hydrolic oil drain out & topped up with


Praveer Sinha 18.05.18 Pranjal B
ity oil in HPU cabinet.
oned by OEM Abhishek Sati 21.05.18

ed by MLA-028 vacum breaker valve inspection


Praveer Sinha 30.05.18
d same is re installed

g done G P Sahoo 21.05.18 Pallab M

s replaced with new one & plate properly


Praveer Sinha 29.05.18 Alton Wr
ok-2 of QRMH-4

sd for replacement after finalization with IOCL

nd modified properly placed. Praveer Sinha 25.05.18 Manoj D

her & o-Ring replaced and leakage arrested Praveer Sinha 31.05.2018

g done Pratap Patra 03.06.18 Rahul Y

ested Praveer Sinha

g done Pratap Patra 30.05.18 Harshit Mishra

g done Pratap Patra 01.06.18 Rahul Y

g done G P Sahoo 18.06.18 Pallab M

al removed from jet/fog nozzele Praveer Sinha 08.06.2018

material removed from TM-04 jet fog movement Praveer Sinha 08.06.18 Pranjal B

x found bend, make shaft & support guide


Praveer Sinha 08.06.18 Pranjal B
p of lathe machine.

g done & flow incation now showing Abhishek Sati 28.06.18 Pallab M

g done & leakage arrested Praveer Sinha 10.06.18 Pranjal B

g done G P Sahoo 06.06.18 Pallab M

g done & made ok Praveer Sinha 08.06.18 Pranjal B


ound in working condition Abhishek Sati 02.07.2018

plate of MD-1 hook-2 replaced with new one Praveer Sinha 18.06.18 B Raju

g done & made ok Praveer Sinha 11.06.18 Rahul Ray

eplaced P Sarvanan 20.06.18 B Raju

nd ok Abhishek Sati 13.6.18 Rahul R

and HOV is free now Praveer Sinha

ound ok Abhishek Sati 27.06.2018

e replaced fpor MLA-04B left right function found


Praveer Sinha 14.06.18 Pranjal B

ound in working condition Abhishek Sati 28.06.2018

nd function can be done with the fixed handle Praveer Sinha

nged Pratap Patra 23.06.2018

th iocl Abhishek Sati 27.06.2018


g done & made ok Abhishek Sati 23.06.2018 Santhosh Reddy

th iocl Abhishek Sati 27.06.2018

ed Praveer Sinha 27.06.2018

made ok after sensor cleaning Abhishek Sati 28.06.2018 Santhosh Reddy

pturehose by new one Praveer Sinha 28.06.18 Raushan Kumar

control room Abhishek Sati 28.06.2018

Praveer Sinha 03.07.18 Raushan Kumar

Field Abhishek Sati 02.07.2018 Rahul Ray

d & spares are under IOCL scope Abhishek Sati 28.06.2018

he end flange Praveer Sinha

g done Pratap Patra 07.07.18 Pranjal B

ng done Pratap Patra 12.07.2018 Pallab M

ield valve is mechanically stuck Abhishek Sati 09.07.18

d in presence of vendor, vendor suggested


Praveer Sinha
PWM card don’t operate FE-103

with new one & leak arrested Praveer Sinha 11.07.18 Pranjal B

CRC, now its ok Pratap Patra 11.07.18 Avdhesh Singh

CRC, now its ok Pratap Patra 11.07.18 Avdhesh Singh

e. Now ok Pratap Patra 12.07.2018 B Raju


nopy fabricated & placed properly Praveer Sinha 08.07.18 Pranjal B

CRC, now its ok Pratap Patra 16.11.2018 Pranjal B

ed for shock dosing inlet & outlet line Praveer Sinha

t available, Follow up under progress with iocl Abhishek Sati 20.07.18

und ok Abhishek Sati 27.7.18

replaced with new one Praveer Sinha 21.07.18 Santhosh Reddy

field Abhishek Sati 26.07.18 Avdhesh Singh

CRC, now its ok Pratap Patra 12.11.2018 Pranjal B

Abhishek Sati 28.07.18

Abhishek Sati 27.07.18 Pranjal B

ulic hose replaced with new hose Praveer Sinha 30.07.18 Ramesh J

ulic hose replaced with new hose Praveer Sinha 06.08.18 Raushan Kumar

o IOCL for Check by OEM. follow up under


Abhishek Sati 29.9.18 Rahul Y

oneywell Abhishek Sati 29.9.18 Rahul Y

ound ok Abhishek Sati 21.08.18 B Raju

heckedusing simulation Abhishek Sati 18.09.18

ound ok Abhishek Sati 21.08.18

hanged with new Pratap Patra 23.08.2018 Ramesh J

ound ok Praveer Sinha 27.08.18 Raushan Kumar

MLA-04A ESD-2 sensor fault rectified Abhishek Sati 13.09.18 Rahul Y


CL-fixed by honeywell rectified loose cabling &
Abhishek Sati 29.09.18
mulation

tified Abhishek Sati 12.09.18 Alton Wr

ted by creating new batch and found ok Abhishek Sati 15.09.18

& Remote mode operation and found ok Abhishek Sati 18.09.18

n performed & now showing working properly Abhishek Sati 18.09.18

Praveer Sinha 25.09.2018 Raushan Kumar

Praveer Sinha 30.09.2018 Raushan Kumar

ce performed. Now working properly. Abhishek Sati 9/24/2018 Santhosh Reddy

Field Abhishek Sati 25.09.18

Praveer Sinha 03.10.2018 Raushan Kumar

Praveer Sinha 05.10.2018 Pallab M

field Checked & found ok Abhishek Sati 21.12.18

broken light Pratap Patra 11.10.2018 Ramesh J

new one Pratap Patra 13.10.18 Santhosh Reddy

reel Praveer Sinha 13.10.18 Manoj D

ermination done Pratap Patra 15.10.18 Pallab M


alled for hose box Praveer Sinha 20.10.18
replaced with new Pratap Patra 16.10.18 Pranjal B
e replaced Praveer Sinha 18.10.18 Santhosh Reddy
ater choked area Pratap Patra 19.10.18 Rahul R
CRC, now its ok Pratap Patra 24.10.18 Pranjal B
field. Remote command not working from LCP,
Abhishek Sati 22.10.18
by Elect Dept.

Pratap patra 22.10.18 Raushan Kumar

tified Abhishek Sati 20.10.18 Pallab M


CRC, now its ok Pratap patra 09.11.2018 Pranjal B
tting struck during operation Abhishek Sati 25.10.18 Pranjal B

Praveer Sinha 25.10.18 Pallab M


Praveer Sinha 30.10.2018 Rahul R
Praveer Sinha 25.10.18 Pallab M
Praveer Sinha 25.10.18 Pallab M
d with IOCL
alled Praveer Sinha 27.10.18 Santhosh Reddy

field Abhishek Sati 31.10.18

ment for TM-4 troubleshooting done Praveer Sinha 29.10.18 Raushan Kumar
es removed Praveer Sinha 29.10.18 Raushan Kumar

CRC, now its ok Pratap Patra 09.11.2018 Pranjal B

CRC, now its ok Pratap Patra 09.11.2018 Pranjal B

Field Abhishek Sati 30.10.18 Alton WR

presence of vendor Praveer Sinha 31.10.18 Rahul R

ked & found ok Praveer Sinha 01.11.18 Rahul R

ling found satisfactory Praveer Sinha Pallab Manna

field Praveer Sinha 03.11.2018 Santhosh Reddy

Praveer Sinha 30.11.2018 Raushan Kumar


ion of JB Pratap Patra 03.11.2018 Rahul R
Praveer Sinha 30.11.2018 Raushan Kumar
Praveer Sinha 30.11.2018 Raushan Kumar

g done Praveer Sinha 10.11.2018 Santhosh Reddy

f IOCL Praveer Sinha

ts Abhishek Sati 27.11.18 Alton WR

found stuck Abhishek Sati 15.11.18


e replaced with new one Praveer Sinha 17.11.18 Rahul R
ontrol room panel Abhishek Sati 28.11.18 Alton WR
ed using welding & M seal Praveer Sinha 04.12.18
and leak arrested Praveer Sinha 04.12.18 Ramesh J

ed using welding & M seal Praveer Sinha 04.12.18

eld Abhishek Sati 26.11.18 Alton WR

ted, checked after calibration & found ok Abhishek Sati 26.12.18

ter Cleaning and Servicing Praveer Sinha 13.12.2018 Santhosh Reddy

eld Abhishek Sati 28.11.18 Alton WR


now ok Pratap Patra 01.12.2018 Raushan Kumar

g of Lighting done. Pratap Patra 06.12.18 Raushan Kumar

hanged Pratap Patra 06.12.2018 Pallab M

valve installed & now it is working properly Praveer Sinha

eed the selector switch Abhishek Sati 17.12.18

eed the selector switch Abhishek Sati 17.12.18

hts fixed out of 12 no's, Remaining will be fixed


Pratap Patra 22.12.18
s to fix now.

ing by tighten all the Gland Praveer Sinha 10.12.2018 Raushan Kumar

ured hose Praveer Sinha 08.12.18 Raushan Kumar

nd ok Abhishek Sati 17.12.18


are fixture used, ok now Pratap Patra 22.12.18

field Abhishek Sati 17.12.2018 Raushan Kumar

done for QCDC cylinder and grease applied Praveer Sinha 13.12.18

new LED Pratap Patra 03.01.19

of the Gate Valve by newly Installed Praveer Sinha 14.12.2018 Raushan Kumar

guard fixed for both condenser pumps Praveer Sinha 20.12.2018

d oiling done Praveer Sinha 17.12.2018 Alton Wr

CV is completed Abhishek Sati 24.12.18 Narasinga Rao B

commissioned alarms Abhishek Sati 26.12.18 Narasinga Rao B

field Abhishek Sati 19.12.2018 B Raju


field Abhishek Sati 24.12.18 Pallab M

ening, cleaning & boxup done P Sarvanan 24.12.18 Ramesh J

P Sarvanan 01.01.19 Rahul R


Pratap Patra 01.01.19 B Raju
f Handle completed Praveer Sinha 07.01.19 Raushan Kumar
Hose replaced Praveer Sinha 04.01.19 Raushan Kumar
new one Praveer Sinha 09.01.19 Rahul R

erminal fitted Pratap Patra 10.01.19

CRC, now its ok Pratap Patra 19.01.19 Narsinga rao

ing by tighten all the flang bolts Praveer Sinha 21.01.19 Narasinga rao B
ing by tighten all the Gland Praveer Sinha 21.01.20 Narasinga rao B

tified card issue from field Abhishek Sati 25.01.19

newnut bolt Pratap Patra 28.01.19

ted Praveer Sinha 21.01.19

ted Praveer Sinha 30.01.19


ted Praveer Sinha 30.01.19
tified card issue from field Abhishek Sati 01.02.19

ing by tighten all the nut bolts Praveer Sinha 09.02.19 Santhosh Reddy
ing done by Mechanical team Praveer Sinha 19.02.19 Rahul y

laced Praveer Sinha 20.04.2019 Santhosh Reddy


tion card loosen from its position due to
SS Malla 30.04.19 Rahul R
was placed on its position & tightened properly

e replaced with new one Praveer Sinha 10.05.19 Ramesh J


Raushan
nesss of Gear,MOV-4305 Working properly SS Malla 10.05.2019
Kumar
Raushan
ration done by Instrument department SS Malla 10.05.2020
Kumar
Raushan
Replaced the swithch by new one Pratap Patra 10.05.2021
Kumar

periodically painting to be done Praveer Sinha 14.05.2019 Rahul Yadav

ration done by Instrument department SS Malla 15.05.2019 Rahul Yadav

maintenance done by Mechanical department Praveer Sinha 15.05.2019 Rahul Yadav


Raushan
ydraulic hose) relaced with new one Praveer Sinha 20.05.2019
Kumar
Rectification done SS Malla 12.06.19 Ramesh J

new & hydraulic oil top up done Gautam Sil 24.06.19 Raushan K

e replaced with new one Bharath R 08.07.19 Ramesh J

tion card loosen from its position due to


K Rana 25.07.2019 K Rana
was placed on its position & tightened properly

e replaced with new one Praveer Sinha 06.08.19 Ramesh J


ed Pratap Patra 26.08.2019 Avdhesh singh

de free Praveer Sinha

cut replaced with new one Ratnakar&Biplab 12.09.2019 Narasinga Rao B

ew one Ramesh J 19.09.2019 Ramesh J


Doc No- IOT/ PBT/ OPNS/ SOJ-LOG / 004

hecked By
REMARKS
Date

02.04.18

09.04.18 Found ok.

07.04.18 Found ok.

07.04.18

18.05.18

10.04.18 No leakage found

10.04.18

12.05.17

10.04.18

19.05.18

13.04.18

15.04.18

16.04.18
25.04.18

03.05.18

18.04.18

22.04.18

26.04.18

04.05.18

07.05.18

08.05.18

10.05.18

11.05.18

13.05.18

12.05.18

18.05.18

18.05.18
21.05.18

29.05.18

25.05.18

03.06.18

30.05.18

01.06.18

19.06.18

08.06.18

08.06.18

19.06.18

11.06.18

07.06.18

08.06.18
19.06.18

No leakage observed during HSD loading to


14.06.18
MT JAG AABHA

20.06.18 Found ok.

Found ok

No Leakage observed after replaced with


14.06.18
new hose
23.06.2018 Found ok.

28.06.2018 Found ok.

28.06.18 found ok

03.07.18

02.07.2018

08.07.18

12.07.2018

11.07.18

11.07.18

11.07.18

12.07.18
08.07.18

16.11.2018 Checked & found OK

21.07.18 Found ok.

26.07.18

16.11.2018 Checked & found OK

27.07.18

30.07.18

06.08.18 Working properly

30.09.18 Working properly

30.09.18 Working properly

21.08.18 Found ok.

23.08.2018 ok

27.08.18 Found ok.

13.09.18 Found ok.


12.09.18 Found ok.

25.09.2018

Hydraullic cabinet canopy remaining

12.10.2018 Checked & found OK

03.10.2018 Checked & found OK

05.10.2018 Checked & found OK

11.10.2018 New one installation is not done

13.10.18 Checked & found OK

13.10.18

16.10.18

17.10.18
18.10.18 Found ok.
19.10.18 Found ok.
16.11.2018 Need to Rust free ladder

To be checkded by electrical department

22.10.2018 Checked during fire engine trail run.

20.10.18 Found ok.


16.11.2018 Checked & found OK
16.11.2018 Checked & found OK
Mechanical Vendor will be done
25.10.18 Found ok.
30.10.18 Checked & Found ok.
25.10.18 Found ok.
25.10.18 Found ok.

27.10.18 Found ok.


For MLA-05B clearance requirement from
Mech Dept.
02.11.18 Found ok.

16.11.2018 Found ok.

16.11.2018 Found ok.

02.11.18 Not working in Remote mode

04.11.18 Found ok.

22.12.18 Found ok.

09.11.18 Checked & found ok

03.11.2018 Found ok.

30.11.2018 Found ok.


17.11.18 Found ok.
30.11.2018 Found ok.
30.11.2018 Found ok.

10.11.2018 HOVs can be operate freely

28.11.18 checked from LCP

17.11.18 Found ok.


28.11.18 checked

06.12.18 Leak observed

28.11.18 checked

13.12.2018 Checked and found ok

28.11.18 Found ok.


03.12.18 Working properly

07.12.18 Good visibility at Pathway

06.12.2018 All tubes are ok and sufficient illumination.

14.12.2018 Founf ok during Vessel Loading

08.12.18 No leaakge during operation


17.12.2018 Shown Properly

14.12.2018 Found ok

18.12.2018 ok

24.12.18

26.12.18

19.12.2018 ok
26.12.18 Checked in DCS & found OK

Leak testing done after boxup & line


24.12.18
charged

01.01.19
01.01.19
07.01.19
04.01.19 Trial Taken after replacemant ,found ok
10.01.19 Checked & found ok

20.01.19 Checked & found ok

21.01.19 Checked & found ok


21.01.19 Checked & found ok

09.02.19 checked & no leakage found


19.02.19 Found ok

21.04.2019 Found ok
Checked & found working properly during
07.05.19 ATF vessel loading to MT DAWN
HARIDWAR

10.05.19

Checked and Found in working condition


10.05.2019
properly
Checked and Found in working condition
10.05.2019
properly

Checked and Found in working condition


10.05.2019
properly

Checked and Found in working condition


14.05.2019
properly
Checked and Found in working condition
15.05.2019
properly
Checked and Found in working condition
15.05.2019
properly
Checked and Found in working condition
20.05.2019
properly
Checked and Found in working condition
12.06.19
properly

24.06.19 Found ok

08.07.19 Found ok

25.07.2019 Found OK

06.08.19 Found OK
26.08.2019

13.09.2019 Found OK

19.09.2019 Found OK

Оценить